Seven-year-old Timothy’s mother takes him to his pediatrician for his annual checkup. His weight is 68 pounds (30.91 kg), plotted at the 95th percentile, and his height is 50 inches (127 cm, 1.27 M), between the 75th and 90th percentiles for his age. Timothy is considered sedentary. His body mass index plots at the 95th percentile for his age. His growth percentiles have been increasing over the last several years. Timothy has a cousin Marley. She is also 7 years old, is considered low active, weighs 69 pound (31.3 Kg) and is 50.5 inches tall (128 cm, 1.28 M). Timothy and Marley have nearly the same Estimated Energy Requirements, 1652 and 1651 Calories per day respectively.

Timothy’s mother expresses concern to the pediatrician about her son’s weight. His older and younger brothers and his cousin Marley are thinner than Timothy. Timothy’s mother is obese, but his father is a normal weight for height. Timothy is in the second grade. He rides the school bus to and from school. He participates in the School Lunch Program at his school. After school, Timothy and his brothers stay in their home with a babysitter until one of their parents returns home from work. Timothy usually watches TV or plays video games after school. His parents leave snack foods—chips, cookies, and sodas—in the house for their sons to have after school. His mother usually prepares their evening meal, which consists of a meat, starch, vegetables, and a dessert item. After dinner, Timothy does his homework and then usually watches more TV with his parents. He usually has a dish of ice cream before going to bed.

  1. Given Timothy and Marley’s age, weight, height, and physical activity level how can their EERs be essentially the same?
  2. What is your assessment of Timothy’s body size based on his weight-for-age, height-for-age, and BMI-for age percentiles?
  3. What suggestions do you have for Timothy’s parents about improving his eating habits?
  4. What suggestions do you have for Timothy’s parents for increasing his physical activity level?
  5. Is it significant that Timothy’s mother also has a weight problem, explain?
Answer #1

Ans

Weight for age = 4× age in years

= 4×7=28 kg

And he has 30.31 kg which is more than expected

Height for age= age×6.5+76(cm)

=7×6.5+76

=45.5+76

=121.5cm

As he has 127 cm height which is also more than expected

BMI=Weight(kg)/height(cm)square

=30.31/1.27×1.27

=18.82BMI is accurate ....he is not under overweight aur obese

Ans

Give small quuantity of food for more number of time and give less trans fat to the child

Ans

For increase physical activity improve the intrest of chiild in outdoor games and parents should have to play with them in starting

Ans

No it is jot sufficient that only her mother is overweight so the child should be overweight because if mother weight gene are dominant then only it lead to the overweight child